# A partially broken Fronsdal model for massless higher-spin particles of integer spin

### v2

@@ 3.2 Spin-3 TDiff model
We have recently suggested [21] a spin-3 analogue of (24), namely,
$\pazocal{L}^{(3)}(a,b,c)=\!-\frac{1}{2}(\partial_{\rho}h_{\mu\nu\alpha})^{2}\!+\!\frac{3}{2}(\partial^{\mu}h_{\mu\nu\alpha})^{2}+3\,a\ \partial^{\mu}\partial^{\nu}h_{\mu\nu\alpha}h^{\alpha}+\frac{3\,b}{2}(\partial_{\rho}h_{\nu})^{2}\!+\!\frac{3\,c}{4}(\partial^{\mu}h_{\mu})^{2}.$ (27)
@@
$\delta h_{\mu\nu\alpha}=\partial_{(\mu}\overline{\Lambda}_{\nu\alpha)}^{T}.$ (28)
The invariants under traceless diffs are given in (18). The transverse condition $\partial^{\nu}\overline{\Lambda}_{\nu\alpha}^{T}=0$ leads, according to (16), to $D$ extra invariants which can be identified with the trace $h_{\mu}$ as in the spin-2 case, since $\delta h_{\mu}=\partial_{\mu}(\eta\cdot\overline{\Lambda}^{T})+2\,\partial^{\nu}\overline{\Lambda}_{\nu\mu}^{T}=0$. The trace $h_{\mu}$ can not be a function of the traceless diff invariants since it is not symmetric under such transformations. Therefore, the trTDiff invariants can be taken as $I_{\tilde{A}}^{trTDiff}=\{I_{A}^{trDiff},h_{\mu}\}$. Similarly to the spin-2 TDiff model, the particle content of (27) can be read off immediately by splitting it into a Fronsdal model plus a trace dependent term via a local field redefinition. Assuming that $a\neq 2/(D+2)$, the required redefinition is unique again:
$h_{\mu\nu\alpha}\rightarrow h_{\mu\nu\alpha}+\frac{1-a}{a(D+2)-2}\eta_{(\mu\nu}h_{\alpha)},$ (29)
@@
$\pazocal{L}^{(3)}(a,b,c)=\pazocal{L}_{F}(I_{ij}^{tt},I_{j}^{t},\tilde{I}_{j}^{t},I,\tilde{I},I_{Tr})+\pazocal{L}_{V}(h_{\mu}),$ (30)
where $\pazocal{L}_{F}(I_{ij}^{tt},I_{j}^{t},\tilde{I}_{j}^{t},I,\tilde{I},I_{Tr})$ is the usual Fronsdal Lagrangian given in (20) and the vector Lagrangian ${\cal{L}}_{V}$ is given only in terms of the trace,
${\cal{L}}_{V}=\frac{3D}{2[a(D+2)-2]^{2}}\left[f_{b}^{(3)}\,h_{\nu}\,\Box\,h^{\nu}+\frac{f_{c}^{(3)}}{2}(\partial^{\mu}h_{\mu})^{2}\right],$ (31)
@@
$f_{b}^{(3)}=D\,(a^{2}-b)+(a-1)^{2}\quad;\quad f_{c}^{(3)}=D\,(c-a^{2})+2\,(a-1)^{2}.$ (32)
Since the trTDiff invariant $h_{\mu}$ can be treated as an independent field coordinate, the particle content of (30) is the content of the usual Fronsdal theory (a physical massless spin-3 particle) plus the content of (31) which is in general a ghostly theory. It is known in massless HS models that the trTDiff symmetry (28) is not enough to get rid of non physical modes and must be enlarged.
If $f^{(3)}_{c}=-2\,f_{b}^{(3)}$, which fine tunes the constant $c$ by
$c=3\,a^{2}-2\,b\equiv c_{s}(a,b),$ (33)
@@ 3.3 The spin-3 PBF model
It is a simple exercise to show that a Lagrangian of the kind $(\partial^{\mu}A_{\mu})^{2}$ has no particle content. On the other hand, from (27), (30) and (31) we see that if we fix $b=b_{S}(a)$ such that $f_{b}^{(3)}=0$, we have a Fronsdal model plus a term proportional to $f_{c}^{(3)}\,(\partial^{\mu}h_{\mu})^{2}$. If $c\neq c_{S}$, which means $f_{c}^{(3)}\,\neq 0$, we stay away from the “Fronsdal family” while still describing only massless spin-3 particles without ghosts. So we are led to define a new massless spin-3 model:
${\cal L}^{(3)}_{PBF}(a,c)\equiv{\cal L}^{(3)}(a,b_{S}(a),c)\quad;\quad c\neq c_{S}(a).$ (35)
@@
${\cal L}^{(3)}_{PBF}\!$ $=$ $\!\!\!-\frac{1}{2}(\partial_{\rho}h_{\mu\nu\alpha})^{2}\!+\!\frac{3}{2}(\partial^{\mu}h_{\mu\nu\alpha})^{2}+3\,\partial^{\mu}\partial^{\nu}h_{\mu\nu\alpha}h^{\alpha}+\frac{3}{2}(\partial_{\rho}h_{\nu})^{2}\!+\!\frac{3c}{4}(\partial^{\mu}h_{\mu})^{2}$ (36)
$=$ ${\cal L}_{F}(I_{\tilde{A}}^{trTDiff})+\frac{3(c-1)}{4}(\partial^{\mu}h_{\mu})^{2},\,$ (37)
@@
$\delta h_{\mu\nu\alpha}=\partial_{(\mu}\bar{\Lambda}^{\ddagger}_{\nu\alpha)}\quad;\quad\eta^{\mu\nu}\bar{\Lambda}^{\ddagger}_{\mu\nu}=0=\partial^{\mu}\partial^{\nu}\bar{\Lambda}^{\ddagger}_{\mu\nu}=0.$ (38)
The extra constraint $\partial^{\mu}\partial^{\nu}\bar{\Lambda}^{\ddagger}_{\mu\nu}=0$ eliminates one gauge parameter and consequently, according to (16), we have one extra gauge invariant besides the traceless diff invariants $I_{A}^{trDiff}$ of the usual Fronsdal model. The extra invariant may be chosen as $\partial^{\mu}h_{\mu}\equiv I_{h}$. Thus, for the PBF model, we can work with the set $I_{\tilde{\tilde{A}}}^{trDiff^{\ddagger}}=\{I_{\tilde{A}}^{trDiff},I_{h}\}$. Next we show why they can be considered as independent field coordinates. First we notice that $I_{h}=\dot{\rho}-\nabla^{2}\Gamma-(D-2)\nabla^{2}\dot{\psi}-\nabla^{2}\dot{\tilde{\psi}}+(D-2)\nabla^{4}\xi+\nabla^{4}\tilde{\xi}$. After the redefinitions $\bar{\Gamma}\equiv\Gamma-\dot{\tilde{\psi}}+\frac{1}{3}\ddot{\tilde{\xi}}$ and $\bar{\tilde{\psi}}\equiv\tilde{\psi}-\frac{\nabla^{2}}{3}\dot{\tilde{\xi}}$, we can write down the matrix taking us from $\{\rho,\Gamma,\psi,\tilde{\psi},\xi,\tilde{\xi}\}$ to $\{I,\bar{\Gamma},\tilde{I},\bar{\tilde{\psi}},I_{tr},I_{h}\}$ as
$\mathbb{M}_{PBF}=\begin{bmatrix}1&-3\partial_{0}&0&3(\partial_{0})^{2}&0&-3(\partial_{0})^{3}\\ 0&1&0&-\partial_{0}&0&\frac{(\partial_{0})^{2}}{3}\\ 0&0&1&0&-\partial_{0}&0\\ 0&0&0&1&0&\frac{-(\partial_{0})}{3}\\ 0&\nabla^{-2}&0&\nabla^{-2}\partial_{0}&D-2&\frac{\Box}{3\nabla^{2}}\\ \nabla^{-4}\partial_{0}&-\nabla^{-2}&-(D-2)\nabla^{-2}\partial_{0}&-\nabla^{-2}\partial_{0}&D-2&1\end{bmatrix}$. (39)
@@ 4.1 Definition
The spin-$s$ analogue of (24) with trTDiff invariance is given by
$\pazocal{L}^{(s)}(a,b,c)=-\frac{1}{2}(\partial h)^{2}+\frac{s}{2}(\partial\!\cdot\!h)^{2}+\frac{a\,s(s-1)}{2}\ \partial\!\cdot\!\partial\!\cdot\!h\ h^{\prime}+\frac{b\,s(s-1)}{4}\ (\partial h^{\prime})^{2}+\frac{c\,s!}{8(s-3)!}\ (\partial\cdot h^{\prime})^{2}$ (40)
for any $a,b,c\in\mathbb{R}$, with $h$ being a rank-s symmetric double-traceless tensor ($h^{{}^{\prime\prime}}=\eta\cdot\eta\cdot h=0$). As in the spin-3 case we implement an invertible r-shift field redefinition in order to split (40) into a usual spin-$s$ Fronsdal model plus extra terms that depend only upon the trace $h^{{}^{\prime}}$. After [^5] The double traceless property is preserved by the r-shift. carrying out in (40):
$h_{\mu_{1}...\mu_{s}}\mapsto h_{\mu_{1}...\mu_{s}}+r\,\eta_{(\mu_{1}\mu_{2}}h^{\alpha}_{\phantom{\alpha}\alpha\mu_{3}...\mu_{s-2})},$ (41)
@@
$r=r(a,s)=\frac{1-a}{a(D+2\,s-4)-2}.$ (43)
Notice that $r(a,s)$ is such that the r-shift is always invertible independently of $(a,s)$ as far as $a\neq a^{*}$. After the r-shift (43) the Lagrangian can be written as the usual Fronsdal model plus two trace dependent terms. Explicitly,
$\pazocal{L}^{(s)}(a,b,c)={\cal L}_{F}^{(s)}(h)+\frac{s(s-1)(D+2s-6)}{4[a(D+2\,s-4)-2]^{2}}\left[f_{b}^{(s)}\,h^{{}^{\prime}}\,\Box\,h^{{}^{\prime}}+\frac{f_{c}^{(s)}(s-2)}{2}\left(\partial\cdot h^{{}^{\prime}}\right)^{2}\right],$ (44)
where ${\cal L}_{F}^{(s)}$ is the usual spin-s Fronsdal Lagrangian corresponding to $\pazocal{L}^{(s)}(1,1,1)$ . We have introduced two constants generalizing [^7] Notice that $f_{D}(a,b)=f_{b}^{(2)}$, see (25). $f_{b}^{(3)}$ and $f_{c}^{(3)}$,
@@
$f_{c}^{(s)}$ $=$ $(D+2s-6)(c-a^{2})+2\,(a-1)^{2}.$ (45)
The corresponding action $S(a,b,c)$ is invariant, in general, under transverse and traceless Diffs (trTDiff). At the “Fronsdal point” $f_{b}^{(s)}=f_{c}^{(s)}=0$ we have a one parameter Fronsdal family invariant under a larger symmetry, i.e., traceless Diff (trDiff). The Fronsdal Lagrangian ${\cal L}_{F}^{(s)}$ must be a function of spin-$s$ trDiff invariants ${\cal L}_{F}^{(s)}={\cal L}_{F}^{(s)}(I_{A}^{trDiff})$. The transverse constraint on the traceless gauge parameters $\partial\cdot\overline{\Lambda}=0$ is a traceless symmetric rank-(s-2) tensor just like the trace $h^{{}^{\prime}}$, which is the natural candidate for the extra invariants raised up by the transverse condition according to (16). So there must exist, for arbitrary integer spin-$s$, a basis of trTDiff invariants given by $I_{\tilde{A}}^{trTDiff}=\{I_{A}^{trDiff},h^{{}^{\prime}}\}$. Since $h^{{}^{\prime}}$ is not invariant under trDiff it must be independent of $I_{A}^{trDiff}$ and the content of (44) can be read off from the Fronsdal model and the last two trace terms independently as in the spin-3 case.
In particular, since $h^{{}^{\prime}}$ is traceless, at the point $f_{c}^{(s)}=2f_{b}^{(s)}$, which requires $c=c_{S}=3\,a^{2}-2\,b$, the $h^{{}^{\prime}}$ terms inside brackets in (44) combine precisely into the irreducible Skvortsov-Vasiliev model [11] for massless spin-(s2)(s-2) particles. So due to the independence of the invariants, the model (44) describes a massless spin-$s$ and a massless spin-(s2)(s-2) particle at $c=c_{S}$. Indeed, we know that at least at the point $c=c_{S}=a=b=0$ the model (40) becomes the rank-s Maxwell-like model of [22] constrained by the double traceless condition. According to [22, 23] it should be a partially reducible model describing massless spin-$s$ and spin-(s2)(s-2) particles, in agreement with our interpretation in terms of independent gauge invariants. This the analogue of the “Maxwell point” of the spin-3 TDiff model of last section. The same spectrum should hold in general for $\pazocal{L}^{(s)}(a,b,c_{S})$ such that $f_{b}^{(s)}\neq 0$.
As in the spin-3 case, besides the “Fronsdal point” and the “TDiff point” $c=c_{S}$, there is another point where we expect (40) to be ghost free, which is our main concern. Namely, at the “PBF point” $f_{b}^{(s)}=0$ with $f_{c}^{(s)}\neq 0$, which requires $c\neq c_{S}$ and
$b=b_{S}^{(s)}=a^{2}+\frac{(a-1)^{2}}{(D+2\,s-6)}.$ (46)
@@
${\cal L}_{PBF}^{(s)}={\pazocal L}^{(s)}(a,b_{S}^{(s)},c)\quad;\quad c\neq 3\,a^{2}-2\,b_{S},$ (47)

### v3

@@ 3.2 Spin-3 TDiff model
We have recently suggested [21] a spin-3 analogue of (3.2), namely,
$\pazocal{L}^{(3)}(a,b,c)=\!-\frac{1}{2}(\partial_{\rho}h_{\mu\nu\alpha})^{2}\!+\!\frac{3}{2}(\partial^{\mu}h_{\mu\nu\alpha})^{2}+3\,a\ \partial^{\mu}\partial^{\nu}h_{\mu\nu\alpha}h^{\alpha}+\frac{3\,b}{2}(\partial_{\rho}h_{\nu})^{2}\!+\!\frac{3\,c}{4}(\partial^{\mu}h_{\mu})^{2}.$ (3.5)
@@
$\delta h_{\mu\nu\alpha}=\partial_{(\mu}\overline{\Lambda}_{\nu\alpha)}^{T}.$ (3.6)
The invariants under traceless diffs are given in (2.18). The transverse condition $\partial^{\nu}\overline{\Lambda}_{\nu\alpha}^{T}=0$ leads, according to (2.16), to $D$ extra invariants which can be identified with the trace $h_{\mu}$ as in the spin-2 case, since $\delta h_{\mu}=\partial_{\mu}(\eta\cdot\overline{\Lambda}^{T})+2\,\partial^{\nu}\overline{\Lambda}_{\nu\mu}^{T}=0$. The trace $h_{\mu}$ can not be a function of the traceless diff invariants since it is not symmetric under such transformations. Therefore, the trTDiff invariants can be taken as $I_{\tilde{A}}^{trTDiff}=\{I_{A}^{trDiff},h_{\mu}\}$. Similarly to the spin-2 TDiff model, the particle content of (3.5) can be read off immediately by splitting it into a Fronsdal model plus a trace dependent term via a local field redefinition. Assuming that $a\neq 2/(D+2)$, the required redefinition is unique again:
$h_{\mu\nu\alpha}\rightarrow h_{\mu\nu\alpha}+\frac{1-a}{a(D+2)-2}\eta_{(\mu\nu}h_{\alpha)},$ (3.7)
@@
$\pazocal{L}^{(3)}(a,b,c)=\pazocal{L}_{F}(I_{ij}^{tt},I_{j}^{t},\tilde{I}_{j}^{t},I,\tilde{I},I_{Tr})+\pazocal{L}_{V}(h_{\mu}),$ (3.8)
where $\pazocal{L}_{F}(I_{ij}^{tt},I_{j}^{t},\tilde{I}_{j}^{t},I,\tilde{I},I_{Tr})$ is the usual Fronsdal Lagrangian given in (2.20) and the vector Lagrangian ${\cal{L}}_{V}$ is given only in terms of the trace,
${\cal{L}}_{V}=\frac{3D}{2[a(D+2)-2]^{2}}\left[f_{b}^{(3)}\,h_{\nu}\,\Box\,h^{\nu}+\frac{f_{c}^{(3)}}{2}(\partial^{\mu}h_{\mu})^{2}\right],$ (3.9)
@@
$f_{b}^{(3)}=D\,(a^{2}-b)+(a-1)^{2}\quad;\quad f_{c}^{(3)}=D\,(c-a^{2})+2\,(a-1)^{2}.$ (3.10)
Since the trTDiff invariant $h_{\mu}$ can be treated as an independent field coordinate, the particle content of (3.8) is the content of the usual Fronsdal theory (a physical massless spin-3 particle) plus the content of (3.9) which is in general a ghostly theory. It is known in massless HS models that the trTDiff symmetry (3.6) is not enough to get rid of non physical modes and must be enlarged.
If $f^{(3)}_{c}=2\,f_{b}^{(3)}$, which fine tunes the constant $c$ by
$c=3\,a^{2}-2\,b\equiv c_{s}(a,b),$ (3.11)
@@ 3.3 The spin-3 PBF model
It is a simple exercise to show that a Lagrangian of the kind $(\partial^{\mu}A_{\mu})^{2}$ has no particle content. On the other hand, from (3.5), (3.8) and (3.9) we see that if we fix $b=b_{S}(a)$ such that $f_{b}^{(3)}=0$, we have a Fronsdal model plus a term proportional to $f_{c}^{(3)}\,(\partial^{\mu}h_{\mu})^{2}$. If $c\neq c_{S}$, which means $f_{c}^{(3)}\,\neq 0$, we stay away from the “Fronsdal family” while still describing only massless spin-3 particles without ghosts. So we are led to define a new massless spin-3 model:
${\cal L}^{(3)}_{PBF}(a,c)\equiv{\cal L}^{(3)}(a,b_{S}(a),c)\quad;\quad c\neq c_{S}(a).$ (3.13)
@@
${\cal L}^{(3)}_{PBF}\!$ $=$ $\!\!\!-\frac{1}{2}(\partial_{\rho}h_{\mu\nu\alpha})^{2}\!+\!\frac{3}{2}(\partial^{\mu}h_{\mu\nu\alpha})^{2}+3\,\partial^{\mu}\partial^{\nu}h_{\mu\nu\alpha}h^{\alpha}+\frac{3}{2}(\partial_{\rho}h_{\nu})^{2}\!+\!\frac{3c}{4}(\partial^{\mu}h_{\mu})^{2}$ (3.14)
$=$ ${\cal L}_{F}(I_{\tilde{A}}^{trTDiff})+\frac{3(c-1)}{4}(\partial^{\mu}h_{\mu})^{2},\,$ (3.15)
@@
$\delta h_{\mu\nu\alpha}=\partial_{(\mu}\bar{\Lambda}^{\ddagger}_{\nu\alpha)}\quad;\quad\eta^{\mu\nu}\bar{\Lambda}^{\ddagger}_{\mu\nu}=0=\partial^{\mu}\partial^{\nu}\bar{\Lambda}^{\ddagger}_{\mu\nu}=0.$ (3.16)
The extra constraint $\partial^{\mu}\partial^{\nu}\bar{\Lambda}^{\ddagger}_{\mu\nu}=0$ eliminates one gauge parameter and consequently, according to (2.16), we have one extra gauge invariant besides the traceless diff invariants $I_{A}^{trDiff}$ of the usual Fronsdal model. The extra invariant may be chosen as $\partial^{\mu}h_{\mu}\equiv I_{h}$. Thus, for the PBF model, we can work with the set $I_{\tilde{\tilde{A}}}^{trDiff^{\ddagger}}=\{I_{\tilde{A}}^{trDiff},I_{h}\}$. Next we show why they can be considered as independent field coordinates. First we notice that $I_{h}=\dot{\rho}-\nabla^{2}\Gamma-(D-2)\nabla^{2}\dot{\psi}-\nabla^{2}\dot{\tilde{\psi}}+(D-2)\nabla^{4}\xi+\nabla^{4}\tilde{\xi}$. After the redefinitions $\bar{\Gamma}\equiv\Gamma-\dot{\tilde{\psi}}+\frac{1}{3}\ddot{\tilde{\xi}}$ and $\bar{\tilde{\psi}}\equiv\tilde{\psi}-\frac{\nabla^{2}}{3}\dot{\tilde{\xi}}$, we can write down the matrix taking us from $\{\rho,\Gamma,\psi,\tilde{\psi},\xi,\tilde{\xi}\}$ to $\{I,\bar{\Gamma},\tilde{I},\bar{\tilde{\psi}},I_{tr},I_{h}\}$ as
$\mathbb{M}_{PBF}=\begin{bmatrix}1&-3\partial_{0}&0&3(\partial_{0})^{2}&0&-3(\partial_{0})^{3}\\ 0&1&0&-\partial_{0}&0&\frac{(\partial_{0})^{2}}{3}\\ 0&0&1&0&-\partial_{0}&0\\ 0&0&0&1&0&\frac{-(\partial_{0})}{3}\\ 0&\nabla^{-2}&0&\nabla^{-2}\partial_{0}&D-2&\frac{\Box}{3\nabla^{2}}\\ \nabla^{-4}\partial_{0}&-\nabla^{-2}&-(D-2)\nabla^{-2}\partial_{0}&-\nabla^{-2}\partial_{0}&D-2&1\end{bmatrix}$. (3.17)
@@ 4.1 Definition
The spin-$s$ analogue of (3.2) with trTDiff invariance is given by
$\pazocal{L}^{(s)}(a,b,c)=-\frac{1}{2}(\partial h)^{2}+\frac{s}{2}(\partial\!\cdot\!h)^{2}+\frac{a\,s(s-1)}{2}\ \partial\!\cdot\!\partial\!\cdot\!h\ h^{\prime}+\frac{b\,s(s-1)}{4}\ (\partial h^{\prime})^{2}+\frac{c\,s!}{8(s-3)!}\ (\partial\cdot h^{\prime})^{2}$ (4.1)
for any $a,b,c\in\mathbb{R}$, with $h$ being a rank-s symmetric double-traceless tensor ($h^{{}^{\prime\prime}}=\eta\cdot\eta\cdot h=0$). As in the spin-3 case we implement an invertible r-shift field redefinition in order to split (4.1) into a usual spin-$s$ Fronsdal model plus extra terms that depend only upon the trace $h^{{}^{\prime}}$. After [^5] The double traceless property is preserved by the r-shift. carrying out in (4.1):
$h_{\mu_{1}...\mu_{s}}\mapsto h_{\mu_{1}...\mu_{s}}+r\,\eta_{(\mu_{1}\mu_{2}}h^{\alpha}_{\phantom{\alpha}\alpha\mu_{3}...\mu_{s-2})},$ (4.2)
@@
$r=r(a,s)=\frac{1-a}{a(D+2\,s-4)-2}.$ (4.4)
Notice that $r(a,s)$ is such that the r-shift is always invertible independently of $(a,s)$ as far as $a\neq a^{*}$. After the r-shift (4.4) the Lagrangian can be written as the usual Fronsdal model plus two trace dependent terms. Explicitly,
$\pazocal{L}^{(s)}(a,b,c)={\cal L}_{F}^{(s)}(h)+\frac{s(s-1)(D+2s-6)}{4[a(D+2\,s-4)-2]^{2}}\left[f_{b}^{(s)}\,h^{{}^{\prime}}\,\Box\,h^{{}^{\prime}}+\frac{f_{c}^{(s)}(s-2)}{2}\left(\partial\cdot h^{{}^{\prime}}\right)^{2}\right],$ (4.5)
where ${\cal L}_{F}^{(s)}$ is the usual spin-s Fronsdal Lagrangian corresponding to $\pazocal{L}^{(s)}(1,1,1)$ . We have introduced two constants generalizing [^7] Notice that $f_{D}(a,b)=f_{b}^{(2)}$, see (3.3). $f_{b}^{(3)}$ and $f_{c}^{(3)}$,
@@
$f_{c}^{(s)}$ $=$ $(D+2s-6)(c-a^{2})+2\,(a-1)^{2}.$ (4.6)
The corresponding action $S(a,b,c)$ is invariant, in general, under transverse and traceless Diffs (trTDiff). At the “Fronsdal point” $f_{b}^{(s)}=f_{c}^{(s)}=0$ we have a one parameter Fronsdal family invariant under a larger symmetry, i.e., traceless Diff (trDiff). The Fronsdal Lagrangian ${\cal L}_{F}^{(s)}$ must be a function of spin-$s$ trDiff invariants ${\cal L}_{F}^{(s)}={\cal L}_{F}^{(s)}(I_{A}^{trDiff})$. The transverse constraint on the traceless gauge parameters $\partial\cdot\overline{\Lambda}=0$ is a traceless symmetric rank-(s-2) tensor just like the trace $h^{{}^{\prime}}$, which is the natural candidate for the extra invariants raised up by the transverse condition according to (2.16). So there must exist, for arbitrary integer spin-$s$, a basis of trTDiff invariants given by $I_{\tilde{A}}^{trTDiff}=\{I_{A}^{trDiff},h^{{}^{\prime}}\}$. Since $h^{{}^{\prime}}$ is not invariant under trDiff it must be independent of $I_{A}^{trDiff}$ and the content of (4.5) can be read off from the Fronsdal model and the last two trace terms independently as in the spin-3 case.
In particular, since $h^{{}^{\prime}}$ is traceless, at the point $f_{c}^{(s)}=2f_{b}^{(s)}$, which requires $c=c_{S}=3\,a^{2}-2\,b$, the $h^{{}^{\prime}}$ terms inside brackets in (4.5) combine precisely into the irreducible Skvortsov-Vasiliev model [11] for massless spin-(s2)(s-2) particles. So due to the independence of the invariants, the model (4.5) describes a massless spin-$s$ and a massless spin-(s2)(s-2) particle at $c=c_{S}$. Indeed, we know that at least at the point $c=c_{S}=a=b=0$ the model (4.1) becomes the rank-s Maxwell-like model of [22] constrained by the double traceless condition. According to [22, 23] it should be a partially reducible model describing massless spin-$s$ and spin-(s2)(s-2) particles, in agreement with our interpretation in terms of independent gauge invariants. This the analogue of the “Maxwell point” of the spin-3 TDiff model of last section. The same spectrum should hold in general for $\pazocal{L}^{(s)}(a,b,c_{S})$ such that $f_{b}^{(s)}\neq 0$.
As in the spin-3 case, besides the “Fronsdal point” and the “TDiff point” $c=c_{S}$, there is another point where we expect (4.1) to be ghost free, which is our main concern. Namely, at the “PBF point” $f_{b}^{(s)}=0$ with $f_{c}^{(s)}\neq 0$, which requires $c\neq c_{S}$ and
$b=b_{S}^{(s)}=a^{2}+\frac{(a-1)^{2}}{(D+2\,s-6)}.$ (4.7)
@@
${\cal L}_{PBF}^{(s)}={\pazocal L}^{(s)}(a,b_{S}^{(s)},c)\quad;\quad c\neq 3\,a^{2}-2\,b_{S},$ (4.8)